# Functions of bounded Musielak-Orlicz-type deformation and anisotropic Total Generalized Variation for image-denoising problems

### v3

@@ 1 Introduction
The extension of the model in [17] to incorporate the case $c=1$, as well as an Orlicz generalization for regularizers $\mathcal{R}(u)\simeq\int_{\Omega}\varphi(x,|Du|)$ for $\Phi$-functions $\varphi$ possibly having linear growth, is the subject of [26], where a thorough mathematical foundation of the space of anisotropic functions with bounded variation with generalized Orlicz growth has also been provided. A further step in this direction has been taken in [28], where the associated Euler-Lagrange equations and gradient flows have also been analyzed. The effect of an anisotropic norm in the Chan-Vese (or anisotropic Mumford-Shah) is the subject of [38]. Anisotropic models, at least in the ROF setting, have indeed proven to allow for more accurate computations, as well as for a staircasing reduction and a sharper texture recognition. We refer to [18, 27, 36] for an overview and further references.
@@ 2.3 Generalized $\Phi$-functions and Orlicz spaces
Thanks to [29, Lemma 3.2.2], it is known that $\|\cdot\|_{\varphi}$ is a quasinorm in $L^{\varphi}(\Omega)$ if $\varphi\in\Phi_{w}(\Omega)$, and a norm if $\varphi\in\Phi_{c}(\Omega)$.
@@
Thanks to [29, Theorem 3.4.6], it is well know that for weak $\Phi$-functions the associate space corresponds to the generalized Orlicz space $L^{\varphi^{*}}(\Omega)$, where
@@
Conditions (i), (ii) and (iv) have been proven in [29], respectively in Lemma 2.4.1, Proposition 2.4.9 and Lemmas 3.7.6 and 4.1.7, while condition (iii) has been shown in [24, Corollary 2.6.3], see also [29, Proposition 2.4.5].
@@
Note that the second continuous inclusion comes from the fact that every weak $\Phi$-function satisfies (aInc)1, see Remark 2.8. This Lemma was shown in [29, Corollary 3.7.9 & Corollary 3.7.10]. We point out that condition (A0) is essential both for establishing the above embeddings and to ensure that $L^{\varphi}(\Omega)$ is a Banach space, cf. [29, Theorem 3.7.13]. In particular, of (A0) is violated, the first inclusion in Lemma 2.13 might fail to hold. We refer to [29, Example 3.7.11] for a counterexample.
@@ 3 The space of bounded deformations fields with generalized Orlicz growth
In the same way, if $\Omega$ is also bounded and $C^{1}$, by [45, Chapter II, Theorem 2.3], having $Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$ implies $u\in BD(\Omega)$. Namely, in the special case $v=Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$, (3.11) rewrites as
@@
We point out that, here and in the following, the set $\Omega$ is only required to have a $C^{1}$ boundary in order to apply [45, Chapter II, Theorem 2.3]. The extension of the above result to the case of Lipschitz boundary will be the subject of the forthcoming paper [23].
@@
This result entails that if $\varphi\in\Phi_{w}(\Omega)$, then $\|\cdot\|_{BD^{\varphi}(\Omega)}$ is a quasinorm. In addition, if $\varphi$ verifies (A0), or if $\varphi\in\Phi_{c}(\Omega)$, then $\|\cdot\|_{BD^{\varphi}(\Omega)}$ is a norm, see [29, Lemma 3.2.2 & Theorem 3.7.13].
@@
differs from the space $BD^{\varphi}(\Omega)$. Thanks to [29, Proposition 2.4.5 & Proposition 3.2.4] it is possible to prove that $BD^{\varphi}(\Omega)=BD^{\varphi^{**}}(\Omega)$, indeed the norms $\|\cdot\|_{\varphi}$ and $\|\cdot\|_{\varphi^{**}}$ are equivalent and the variations coincide, implying that the norms $\|\cdot\|_{BD^{\varphi}(\Omega)}$ and $\|\cdot\|_{BD^{\varphi^{**}}(\Omega)}$ are equivalent as well. Note that some results needed for this consideration, e.g. [29, Lemma 2.4.3] or [29, Proposition 2.4.5], are only stated without the $x$-dependence, but they hold true even in the generalized Orlicz setting, while others, e.g. [29, Theorem 2.5.10] or [29, Proposition 3.2.4], are already stated with the $x$-dependence.
With the next example we want to show that, at least considering some $\Phi$-functions for which a Korn-type inequality does not hold, the space $BV^{\varphi}(\Omega)$ is a proper subset of $BD^{\varphi}(\Omega)$. The example follows the lines of [3, Example 7.7], exploiting a construction developed in [41], see also [45, Chapter I, Subsection 1.3].
@@
Note that many $\Phi$-functions, e.g. many out of the ones listed in [29, Example 2.5.3], fulfill the assumption of coinciding with $|t|$ on a ball $B$, meaning that this example holds true in many contexts.
The validity of Korn-type inequalities in the context of Orlicz spaces has been studied in different contexts, for example [14, Theorem 1.1] where it was shown that the $\nabla_{2}$-condition and the $\Delta_{2}$-condition are sharp conditions for Korn-type inequalities in Orlicz spaces (for the definition of these two conditions see [42, Subsection 2.3] or [13, Equations (2.6) & (2.7)]). In [20, Theorems 3.1 & 3.3] the validity of Korn-type inequalities for Young functions was shown, while in [13, Theorems 3.1 & 3.3] the equivalence of these Korn-type inequalities to two balance conditions, similar to the $\nabla_{2}$- and $\Delta_{2}$-conditions, has been proven. Note that, under the assumptions considered in [20] and [13], their definition of Young functions and our definition of convex $\Phi$-functions, when independent of $x$, are equivalent. In the context of generalized Orlicz spaces (or Musielak-Orlicz spaces) instead, there are results in the case of variable exponents Lebesgue spaces when the exponent is bounded away from $1$, see [25, Section 5] and [24, Theorems 14.3.21 & 14.3.23], but not much else is known.
@@ 3.1 Weighted modular and associated norms
On the one hand, note that testing (3.15) with $\psi=0$, we get that the supremum in $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}$ is always non-negative. On the other hand, if $\varrho_{\varphi^{*}}(|\psi|)=+\infty$, then the integral in (3.15) is equal to -$\infty$, since the first term is finite. Here we used again [45, Chapter II, Theorem 2.3] yielding that $Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$ implies $u\in BD(\Omega)$.
@@
entailing that for any $\lambda<1$ it holds $\|v\|_{\varrho_{\widetilde{V}^{n\times n}_{\varphi}}}>\lambda$. By [29, Lemma 3.2.3] then this implies that $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}(\tfrac{v}{\lambda})>1$ for any $\lambda<1$. On the other hand, having $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}(\tfrac{v}{\lambda})>1$ for any $\lambda<1$ entails that, by Definition 2.3,
@@
directly from the definition of $\inf$. Now, by [29, Lemma 3.2.3] and thanks to the fact that $\varphi^{*}\in\Phi_{c}(\Omega)$, see Lemma 2.12(iii), $\varrho_{\varphi^{*}}(|w|)\leq 1$ if and only if $\|w\|_{\varphi^{*}}\leq 1$, and we can immediately conclude that
@@
In the second part of Lemma 3.13 we assume that $C^{1}_{c}(\Omega;\mathbb{R}^{n\times n})$ is dense in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{n\times n})$. If $\varphi^{*}$ satisfies (A0) and (aDec), then this holds automatically by [29, Theorem 3.7.15] (note that this theorem is stated for scalar functions, but the vectorial case can be deduced working components-wise thanks to the equivalence of euclidean norms in spaces of finite dimension). We know already that, by Lemma 2.12(ii) & (iv), $\varphi^{*}$ satisfies these conditions if and only if $\varphi$ satisfies (A0) and (aInc), meaning that it is not clear whether this density is there or not in the subsets where the function $\varphi$ is linear.
@@ 3.2 Explicit expression of the anisotropic modular
In [30, Section 3 & Example A.1] it was shown that log-Hölder continuity is not sufficient when working in $BV^{p(\cdot)}$ and here, similarly, the (A1) condition (corresponding to log-Hölder continuity in the general case of convex $\Phi$-functions) is not sufficient in the previous result, as in the following ones, in view of [26, Example 4.3]. That is why we make use of the restricted (VA1) condition, which corresponds to strong log-Hölder continuity in the case of the variable exponent function, see [26, Proposition 3.2].
@@ 3.3 The variable exponent case
with $p:\Omega\to[1,+\infty)$ a measurable function. Let us recall that, choosing $\varphi$ as in (3.22) and considering the set
@@
respectively $q^{\pm}_{A}$ for the conjugate exponent function $q$. In the following, we will assume $p^{+}_{\Omega}$$\infty$.
@@
The first result we prove in this subsection is the equivalent of Theorem 3.14 in this specific setting. In particular, we show that the same conclusion of Theorem 3.14 holds under less restrictive hypotheses on our function $\varphi$. Before doing so, we need to recall the definition of the modular associated to the variable exponent space when $p^{+}_{\Omega}$$\infty$, given in an equivalent way to the one in [30, Subsection 2.2], namely
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded, connected and open set. Let $\varphi$ be defined as in (3.22) and let $p:\Omega\to[1,+\infty)$ be a lower semicontinuos function with $p^{+}_{\Omega}$$\infty$ and such that $Y$ is closed, see (3.23). If $u\in BD(\Omega)$, then
@@
First, let us observe that the claim of Lemma 3.16 is automatically satisfied in the case of $\varphi$ given by (3.22) with no continuity required on $p$, but only under the assumption that $p^{+}_{\Omega}$ is finite. Indeed, on the one hand, in view of (3.24) and Definition 2.9 (which is a limit in this specific case), it suffices to prove the inequality in the set $Y$ given by (3.23), since in $\Omega\setminus Y$ the recession function is $\infty$. On the other hand though, whenever $\psi\in C(\Omega)$ with $\varrho_{\varphi^{*}}(|\psi|)<+\infty$, then by (3.24) we have that $\varphi^{*}(x,|\psi|)=0$ for almost every $x\in Y$, implying that $|\psi|\leq 1$ almost everywhere in $Y$. This means that $|\psi(x)|\leq 1=\varphi^{\infty}(x)$ in $Y$, which proves our claim.
As a second step, let us remark that choosing $\varphi$ as in (3.22), a careful inspection of the proof of Proposition 3.18 guarantees that the result is still true if we remove the assumptions therein and, again, we only require that $p^{+}_{\Omega}$$\infty$. In fact, by Remark 3.19, we know that (A0) and (aDec) are satisfied and this implies that we can apply [29, Theorem 3.7.15], meaning the required continuity on $\varphi$ can be bypassed exploiting the density of $C^{\infty}_{c}(\Omega)$ in $L^{\varphi}(\Omega)$. Hence, we can conclude that
@@
Thanks to the lower semicontinuity assumption on $p$ on $\Omega$, we obtain that (A.1) and (A.2) hold. In particular, the second one holds for any $\beta\in[0,1]$, due to the non decreasing monotonicity of $(\cdot)^{p(x)}$. Observe that $h_{k}$ is lower semicontinuous by definition (it is a lower semicontinuous envelope, see [21, Chapter 1]) and increasing in $k$. Consequently the same arguments as in the proof of Proposition 3.17 lead to the desired conclusion. Moreover, let us remark that for every $u\in BD(\Omega)$ it holds
@@
when $\varphi$ is as in (3.22). Alternatively, this latter equality can be proven directly arguing as in [8, Theorem 3.8].

### v4

@@ 1 Introduction
The extension of the model in [17] to incorporate the case $c=1$, as well as an Orlicz generalization for regularizers $\mathcal{R}(u)\simeq\int_{\Omega}\varphi(x,|Du|)$ for $\Phi$-functions $\varphi$ possibly having linear growth, is the subject of [26], where a thorough mathematical foundation of the space of anisotropic functions with bounded variation with generalized Orlicz growth has also been provided. A further step in this direction has been taken in [29], where the associated Euler-Lagrange equations and gradient flows have also been analyzed (see also [28]). The effect of an anisotropic norm in the Chan-Vese (or anisotropic Mumford-Shah) is the subject of [39]. Anisotropic models, at least in the ROF setting, have indeed proven to allow for more accurate computations, as well as for a staircasing reduction and a sharper texture recognition. We refer to [18, 27, 37] for an overview and further references.
@@ 2.3 Generalized $\Phi$-functions and Orlicz spaces
Thanks to [30, Lemma 3.2.2], it is known that $\|\cdot\|_{\varphi}$ is a quasinorm in $L^{\varphi}(\Omega)$ if $\varphi\in\Phi_{w}(\Omega)$, and a norm if $\varphi\in\Phi_{c}(\Omega)$.
@@
Thanks to [30, Theorem 3.4.6], it is well know that for weak $\Phi$-functions the associate space corresponds to the generalized Orlicz space $L^{\varphi^{*}}(\Omega)$, where
@@
Conditions (i), (ii) and (iv) have been proven in [30], respectively in Lemma 2.4.1, Proposition 2.4.9 and Lemmas 3.7.6 and 4.1.7, while condition (iii) has been shown in [24, Corollary 2.6.3], see also [30, Proposition 2.4.5].
@@
Note that the second continuous inclusion comes from the fact that every weak $\Phi$-function satisfies (aInc)1, see Remark 2.8. This Lemma was shown in [30, Corollary 3.7.9 & Corollary 3.7.10]. We point out that condition (A0) is essential both for establishing the above embeddings and to ensure that $L^{\varphi}(\Omega)$ is a Banach space, cf. [30, Theorem 3.7.13]. In particular, of (A0) is violated, the first inclusion in Lemma 2.13 might fail to hold. We refer to [30, Example 3.7.11] for a counterexample.
@@ 3 The space of bounded deformations fields with generalized Orlicz growth
In the same way, if $\Omega$ is also bounded and $C^{1}$, by [46, Chapter II, Theorem 2.3], having $Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$ implies $u\in BD(\Omega)$. Namely, in the special case $v=Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$, (3.11) rewrites as
@@
We point out that, here and in the following, the set $\Omega$ is only required to have a $C^{1}$ boundary in order to apply [46, Chapter II, Theorem 2.3]. The extension of the above result to the case of Lipschitz boundary will be the subject of the forthcoming paper [23].
@@
This result entails that if $\varphi\in\Phi_{w}(\Omega)$, then $\|\cdot\|_{BD^{\varphi}(\Omega)}$ is a quasinorm. In addition, if $\varphi$ verifies (A0), or if $\varphi\in\Phi_{c}(\Omega)$, then $\|\cdot\|_{BD^{\varphi}(\Omega)}$ is a norm, see [30, Lemma 3.2.2 & Theorem 3.7.13].
@@
differs from the space $BD^{\varphi}(\Omega)$. Thanks to [30, Proposition 2.4.5 & Proposition 3.2.4] it is possible to prove that $BD^{\varphi}(\Omega)=BD^{\varphi^{**}}(\Omega)$, indeed the norms $\|\cdot\|_{\varphi}$ and $\|\cdot\|_{\varphi^{**}}$ are equivalent and the variations coincide, implying that the norms $\|\cdot\|_{BD^{\varphi}(\Omega)}$ and $\|\cdot\|_{BD^{\varphi^{**}}(\Omega)}$ are equivalent as well. Note that some results needed for this consideration, e.g. [30, Lemma 2.4.3] or [30, Proposition 2.4.5], are only stated without the $x$-dependence, but they hold true even in the generalized Orlicz setting, while others, e.g. [30, Theorem 2.5.10] or [30, Proposition 3.2.4], are already stated with the $x$-dependence.
With the next example we want to show that, at least considering some $\Phi$-functions for which a Korn-type inequality does not hold, the space $BV^{\varphi}(\Omega)$ is a proper subset of $BD^{\varphi}(\Omega)$. The example follows the lines of [3, Example 7.7], exploiting a construction developed in [42], see also [46, Chapter I, Subsection 1.3].
@@
Note that many $\Phi$-functions, e.g. many out of the ones listed in [30, Example 2.5.3], fulfill the assumption of coinciding with $|t|$ on a ball $B$, meaning that this example holds true in many contexts.
The validity of Korn-type inequalities in the context of Orlicz spaces has been studied in different contexts, for example [14, Theorem 1.1] where it was shown that the $\nabla_{2}$-condition and the $\Delta_{2}$-condition are sharp conditions for Korn-type inequalities in Orlicz spaces (for the definition of these two conditions see [43, Subsection 2.3] or [13, Equations (2.6) & (2.7)]). In [20, Theorems 3.1 & 3.3] the validity of Korn-type inequalities for Young functions was shown, while in [13, Theorems 3.1 & 3.3] the equivalence of these Korn-type inequalities to two balance conditions, similar to the $\nabla_{2}$- and $\Delta_{2}$-conditions, has been proven. Note that, under the assumptions considered in [20] and [13], their definition of Young functions and our definition of convex $\Phi$-functions, when independent of $x$, are equivalent. In the context of generalized Orlicz spaces (or Musielak-Orlicz spaces) instead, there are results in the case of variable exponents Lebesgue spaces when the exponent is bounded away from $1$, see [25, Section 5] and [24, Theorems 14.3.21 & 14.3.23], but not much else is known.
@@ 3.1 Weighted modular and associated norms
On the one hand, note that testing (3.15) with $\psi=0$, we get that the supremum in $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}$ is always non-negative. On the other hand, if $\varrho_{\varphi^{*}}(|\psi|)=+\infty$, then the integral in (3.15) is equal to -$\infty$, since the first term is finite. Here we used again [46, Chapter II, Theorem 2.3] yielding that $Eu\in\mathcal{M}(\Omega;\mathbb{R}^{n\times n}_{\rm sym})$ implies $u\in BD(\Omega)$.
@@
entailing that for any $\lambda<1$ it holds $\|v\|_{\varrho_{\widetilde{V}^{n\times n}_{\varphi}}}>\lambda$. By [30, Lemma 3.2.3] then this implies that $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}(\tfrac{v}{\lambda})>1$ for any $\lambda<1$. On the other hand, having $\varrho_{\widetilde{V}^{n\times n}_{\varphi}}(\tfrac{v}{\lambda})>1$ for any $\lambda<1$ entails that, by Definition 2.3,
@@
directly from the definition of $\inf$. Now, by [30, Lemma 3.2.3] and thanks to the fact that $\varphi^{*}\in\Phi_{c}(\Omega)$, see Lemma 2.12(iii), $\varrho_{\varphi^{*}}(|w|)\leq 1$ if and only if $\|w\|_{\varphi^{*}}\leq 1$, and we can immediately conclude that
@@
In the second part of Lemma 3.13 we assume that $C^{1}_{c}(\Omega;\mathbb{R}^{n\times n})$ is dense in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{n\times n})$. If $\varphi^{*}$ satisfies (A0) and (aDec), then this holds automatically by [30, Theorem 3.7.15] (note that this theorem is stated for scalar functions, but the vectorial case can be deduced working components-wise thanks to the equivalence of euclidean norms in spaces of finite dimension). We know already that, by Lemma 2.12(ii) & (iv), $\varphi^{*}$ satisfies these conditions if and only if $\varphi$ satisfies (A0) and (aInc), meaning that it is not clear whether this density is there or not in the subsets where the function $\varphi$ is linear.
@@ 3.2 Explicit expression of the anisotropic modular
In [31, Section 3 & Example A.1] it was shown that log-Hölder continuity is not sufficient when working in $BV^{p(\cdot)}$ and here, similarly, the (A1) condition (corresponding to log-Hölder continuity in the general case of convex $\Phi$-functions) is not sufficient in the previous result, as in the following ones, in view of [26, Example 4.3]. That is why we make use of the restricted (VA1) condition, which corresponds to strong log-Hölder continuity in the case of the variable exponent function, see [26, Proposition 3.2].
@@ 3.3 The variable exponent case
with $p:\Omega\to[1,+\infty)$ a measurable function. Such a $p$ is called a variable exponent. Let us recall that, choosing $\varphi$ as in (3.22) and considering the set
@@
respectively $q^{\pm}_{A}$ for the conjugate exponent function $q$. In the following, we will assume $p^{+}_{\Omega}$$\infty$. We start by recalling the following notion, introduced in [8].
Definition 3.19.
A variable exponent $p$ is strongly $\log$-Hölder continuous in $Y$ if
$p(x)-1\leq\frac{\omega(|x-y|)}{\log(e+\frac{1}{|x-y|})},$
for some $\omega:[0,\infty)\to[0,\infty)$ with $\lim_{t\to 0}\omega(t)=0$ and every $y\in Y$ and $x\in\Omega$.
At this point, it is worth observing that the above property is weaker than the ’classical $\log$-Hölder continuity’ as stated in [8, Definition 2.3], which is, in turn, equivalent to the restricted (VA1) condition, introduced in Definition 2.10, for the function (3.22).
In [8] the following technical result has been proven.
Lemma 3.20.
Suppose that $p$ is strongly $\log$-Hölder continuous in $Y$. If $\rho_{q(\cdot)}(w)<\infty$ for some continuous $w$, then $|w|\leq 1$ in $Y$.
@@
The first result we prove in this subsection is the equivalent of Theorem 3.14 in this specific setting. In particular, we show that the same conclusion of Theorem 3.14 holds under less restrictive hypotheses on our function $\varphi$. Before doing so, we need to recall the definition of the modular associated to the variable exponent space when $p^{+}_{\Omega}$$\infty$, given in an equivalent way to the one in [31, Subsection 2.2], namely
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded, connected and open set. Let $\varphi$ be defined as in (3.22) and let $p:\Omega\to[1,+\infty)$ be continuous in $\Omega$ and strongly $\log$-Hölder continuous in $Y$, with $p^{+}_{\Omega}$$\infty$. If $u\in BD(\Omega)$, then
@@
First, let us observe that the claim of Lemma 3.16 is automatically satisfied in the case of $\varphi$ given by (3.22) with the only requirement of strong $\log$-Hölder continuity in $Y$ and the assumption $p^{+}_{\Omega}$ is finite. Indeed, on the one hand, in view of (3.24) and Definition 2.9 (which is a limit in this specific case), it suffices to prove the inequality in the set $Y$ given by (3.23), since in $\Omega\setminus Y$ the recession function is $\infty$. On the other hand though, whenever $\psi\in C(\Omega)$ with $\varrho_{\varphi^{*}}(|\psi|)<+\infty$, then by (3.24) we have that $\varphi^{*}(x,|\psi|)=0$ for almost every $x\in Y$, implying that $|\psi|\leq 1=\varphi^{\infty}(x)$ in $Y$, in view of Lemma 3.20, which proves our claim.
As a second step, let us remark that choosing $\varphi$ as in (3.22), with $p(\cdot)$ continuous, , Proposition 3.18 is valid. In fact, by assumption $\varphi$ is continuous, and by Remark 3.21, we know that (A0) and (aDec) are satisfied. Hence, we can conclude that
@@
Thanks to the continuity assumption on $p$ in $\Omega$, we obtain that (A.1) and (A.2) hold. In particular, the second one holds for any $\beta\in[0,1]$, due to the non decreasing monotonicity of $(\cdot)^{p(x)}$. Observe that $h_{k}$ is lower semicontinuous by definition (it is a lower semicontinuous envelope, see [21, Chapter 1]) and increasing in $k$. Consequently the same arguments as in the proof of Proposition 3.17 lead to the desired conclusion. Moreover, let us remark that for every $u\in BD(\Omega)$ it holds
@@
when $\varphi$ is as in (3.22). Alternatively, this latter equality can be proven directly arguing as in [8, Theorem 3.8] even removing the continuity hypotesis on $p(\cdot)$, replacing it by lower semicontinuity .